Mr Greycloud works in the lab at a pharmaceutical company. He needs to make 27 liters of a 25% acid soluction to test a new product. His supplier only ships a 44% and a 17% solution. Mr Greycloud decides to make the 25% solution by mixing the 44% solution with the 17% solution. How much of the 44% solution will Mr Greycloud need to use?
Answer by stanbon(75887)

Mr Greycloud works in the lab at a pharmaceutical company. He needs to make 27 liters of a 25% acid soluction to test a new product. His supplier only ships a 44% and a 17% solution. Mr Greycloud decides to make the 25% solution by mixing the 44% solution with the 17% solution. How much of the 44% solution will Mr Greycloud need to use?
-------
Quantity Eq:: f + s = 27 liters
active Eq:: 44f + 17s = 25*27
--------------------------------------
Modify to solve for "f"::
17f + 17s = 17*27
44f + 17s = 25*27
-------
Subtract and solve for "f":
27f = 8*27
f = 8 liters (amt. of 44% solution needed)
--------------------------------------------------
Cheers,
Stan H.
----------------------